OpenStudy (anonymous):

If the width of a rectangle is x + 2 and the length is twice as long, what is the area? A. 6x + 12 B. 2x² + 6x + 4 C. 2x² + 8x + 6 D. 2x² + 8x + 8

OpenStudy (anonymous):

width , x+2 length , 2(x+2) area = length x width =2(x+2)(x+2) = 2(x+2)^2 =2x^2+8x +8

OpenStudy (anonymous):

(x+2)(2x+4) 2x^2+4x+4x+8 2x^2+8x++8